OpenStudy (darkprince14):

Integrate: cosx/ sqrt(1-sinx) from 0 to pi/2

hartnn (hartnn):

did u try to put u = 1-sin x ??

hartnn (hartnn):

then du =... ?

hartnn (hartnn):

if you are doing this (substitution) way, don't forget to change the limits when x= 0, u= ..... ? when x=pi/2, u= ... ?

hartnn (hartnn):

@darkprince14 still here ?

OpenStudy (darkprince14):

yep..sorry.. i was doing other math problems..

hartnn (hartnn):

ok, so you want to try my method ?

OpenStudy (darkprince14):

Isn't it an improper integral?

OpenStudy (darkprince14):

the answer is 2 using the method you gave me.. is this correct?

hartnn (hartnn):

yup! exactly correct, its 2.

OpenStudy (darkprince14):

what would the answer be if we use the rules on improper integrals? I tried integrating it but i just can't get it.. we have a Long Exam tomorrow... any tips for me?

OpenStudy (darkprince14):

Thank you for guiding me @hartnn

hartnn (hartnn):

you having trouble with improper integrals? any particular integral you have trouble with, which use the rules? this one doesn't...

OpenStudy (darkprince14):

\[\int\limits_{0}^{+\infty} \frac{ 3^{-\sqrt{x}}dx }{ \sqrt{x} }\] this one I can't get...

OpenStudy (darkprince14):

@hartnn

terenzreignz (terenzreignz):

Well, let's first focus on getting the antiderivative... \[\Large \int \frac{3^{-\sqrt x}}{\sqrt x}dx\]

terenzreignz (terenzreignz):

Any idea what substitution to use? :)

OpenStudy (darkprince14):

-sqrt(x) = u ...

terenzreignz (terenzreignz):

Let's try that :) \[\Large u = -\sqrt x\]\[\Large du = -\frac1{2\sqrt x}dx\]\[\Large dx = -2\sqrt x du\] I assume you got up to this point? :)

OpenStudy (darkprince14):

yep... so if we transfer it to the original equation it'll be \[\int\limits_{0}^{+\infty} 3^{u} du\]

terenzreignz (terenzreignz):

Not quite..., not so fast ^_^ Forget about the limits for now, and just focus on the antiderivative...

terenzreignz (terenzreignz):

\[\Large \int \frac{3^{-\sqrt x}}{\sqrt x}dx\] \[\Large u = -\sqrt x\] \[\Large dx = -2\sqrt x du\]

terenzreignz (terenzreignz):

\[\Large \int \frac{3^{\color{red}u}}{\sqrt x}dx\]\[\Large \int \frac{3^{u}}{\sqrt x}\color{red}{\left(-2\sqrt x du\right)}\]\[\Large -2\int 3^u du\]

OpenStudy (darkprince14):

ooh.. then it's antiderivative will be -2 (3^u/ ln3)

OpenStudy (darkprince14):

\[-2(\frac{ 3^{u} }{ \ln 3 })\]

terenzreignz (terenzreignz):

That is correct :) And bringing it back in terms of x, you get... \[\Large -\frac{2}{\ln 3}\cdot 3^{-\sqrt x}\]

terenzreignz (terenzreignz):

So, It is time to evaluate it from 0 to a positive number, say k...\[\Large\int\limits_{0}^{+\infty} \frac{ 3^{-\sqrt{x}} }{ \sqrt{x} }dx=\lim_{k\rightarrow +\infty}\left[-\frac{2}{\ln 3}\cdot3^{-\sqrt x}\right]_0^{k} \] I trust you got it from here? ^_^

OpenStudy (darkprince14):

yep... may i ask something?it's wrong to multiply 2 to 3^-sqrt(x)

hartnn (hartnn):

sorry, i lost internet connection due to heavy rains :(

OpenStudy (darkprince14):

??? right???

terenzreignz (terenzreignz):

The earth be goin' crazy :3

OpenStudy (darkprince14):

@hartnn you're back^_^ I hope the rain in your place won't be long...

terenzreignz (terenzreignz):

Unless they have the exact same exponent, you do not simply multiply 2 to the \(\Large 3^{-\sqrt x}\)

terenzreignz (terenzreignz):

Compare... \[\Large 2 \cdot 3 ^m = 2\cdot 3^m\](can't be simplified further...) \[\Large 2^m\cdot 3^m = (2\cdot 3)^m=6^m\]

terenzreignz (terenzreignz):

Get it? Got it? Good :) --------------------------------- Terence out